# CLARE: Cognitive Load Assessment in REaltime with Multimodal Data

### v1

@@ VI-A Classification Algorithms
GB LGBM LDA LR MLP RF SVM XGBoost CNN
ECG 66.58 (65.49) 66.71 (66.27) 60.49 (57.62) 60.73 (58.21) 62.19 (61.07) 68.28 (67.38) 64.20 (63.83) 66.68 (65.26) 78.45 (76.41)
EDA 68.34 (67.39) 70.18 (69.64) 63.52 (61.65) 63.48 (61.53) 60.56 (59.03) 71.98 (70.93) 65.59 (65.00) 69.33 (68.06) 68.01 (66.21)
EEG 69.19 (68.36) 69.94 (69.42) 64.30 (63.02) 64.84 (63.62) 62.53 (61.94) 71.85 (71.02) 69.09 (68.67) 69.91 (69.00) 74.10 (72.08)
Gaze 66.85 (66.01) 67.97 (67.45) 62.39 (60.11) 62.50 (60.35) 61.54 (59.61) 69.70 (68.66) 65.83 (65.51) 67.80 (66.72) 66.78 (64.23)
ECG, EDA 70.08 (69.31) 71.98 (71.56) 64.37 (62.89) 64.67 (63.20) 59.95 (58.15) 73.24 (72.51) 69.30 (68.83) 70.96 (70.15) 79.72 (77.53)
ECG, EEG 70.76 (69.99) 70.18 (69.67) 65.66 (64.59) 65.93 (64.96) 63.38 (62.56) 72.53 (71.77) 68.75 (68.33) 69.30 (68.32) 77.66 (75.77)
ECG, Gaze 68.96 (68.11) 70.38 (69.92) 63.96 (62.71) 64.13 (62.92) 63.86 (62.67) 71.00 (70.06) 67.80 (67.33) 70.59 (69.71) 78.41 (76.54)
EDA, EEG 72.19 (71.39) 72.09 (71.63) 65.35 (64.21) 65.28 (64.22) 60.96 (59.17) 73.61 (72.85) 68.86 (68.40) 72.49 (71.65) 75.30 (73.45)
EDA, Gaze 71.34 (70.63) 72.32 (71.87) 65.45 (64.12) 65.73 (64.43) 57.67 (54.13) 74.40 (73.65) 68.85 (68.34) 72.63 (71.83) 70.64 (68.23)
EEG, Gaze 71.64 (70.92) 71.71 (71.21) 66.98 (66.08) 67.43 (66.48) 66.30 (64.85) 73.65 (72.76) 70.79 (70.40) 71.40 (70.59) 75.50 (73.13)
ECG, EDA, EEG 71.10 (70.42) 72.90 (72.57) 66.13 (65.07) 65.73 (64.74) 61.71 (60.51) 73.65 (72.97) 69.87 (69.43) 72.53 (71.72) 80.10 (78.49)
ECG, EDA, Gaze 72.12 (71.45) 73.24 (72.86) 67.12 (66.04) 66.61 (65.60) 62.39 (60.10) 74.53 (73.85) 70.96 (70.49) 73.27 (72.54) 80.30 (78.34)
ECG, EEG, Gaze 71.51 (70.82) 72.80 (72.34) 67.56 (66.74) 68.51 (67.67) 66.54 (65.14) 74.09 (73.26) 71.13 (70.80) 71.64 (70.85) 78.26 (76.56)
EDA, EEG, Gaze 73.58 (72.96) 73.58 (73.13) 67.97 (67.13) 67.87 (67.03) 61.51 (59.51) 75.18 (74.45) 71.71 (71.26) 74.50 (73.78) 75.87 (74.08)
All modalities 73.38 (72.81) 74.26 (73.83) 68.24 (67.47) 68.38 (67.59) 62.36 (60.66) 75.14 (74.47) 72.05 (71.59) 74.26 (73.60) 79.65 (77.94)
@@
GB LGBM LDA LR MLP RF SVM XGBoost CNN
ECG 53.23 (45.21) 42.00 (33.06) 55.60 (43.79) 43.34 (35.22) 43.15 (39.08) 55.60 (44.16) 40.26 (33.09) 40.00 (28.79) 63.28 (51.00)
EDA 58.67 (50.86) 55.20 (51.03) 60.15 (51.23) 50.81 (47.56) 48.90 (43.91) 59.36 (49.50) 50.46 (47.42) 51.52 (47.67) 64.08 (54.80)
EEG 54.71 (47.06) 53.35 (46.40) 49.76 (43.83) 52.38 (44.16) 54.46 (48.94) 59.17 (46.57) 53.23 (47.03) 56.54 (48.95) 65.21 (52.15)
Gaze 54.61 (47.35) 54.79 (46.41) 56.72 (42.49) 52.79 (43.97) 53.08 (44.76) 58.80 (48.83) 51.63 (44.05) 56.93 (46.53) 60.23 (52.92)
ECG, EDA 58.66 (50.54) 45.88 (38.04) 58.37 (47.25) 51.87 (48.68) 47.60 (44.53) 62.23 (51.60) 50.58 (47.58) 47.50 (43.31) 64.48 (55.38)
ECG, EEG 55.10 (48.49) 53.54 (46.77) 50.27 (44.37) 51.97 (44.11) 52.17 (47.11) 56.32 (43.89) 49.29 (43.36) 56.97 (50.15) 65.22 (51.00)
ECG, Gaze 54.29 (49.07) 56.62 (50.75) 57.21 (47.22) 57.22 (47.42) 53.65 (48.40) 57.66 (47.21) 54.51 (47.36) 56.59 (49.79) 59.27 (50.81)
EDA, EEG 56.53 (49.60) 57.91 (52.19) 51.51 (44.89) 51.27 (44.51) 52.80 (46.45) 60.33 (48.58) 53.19 (47.38) 62.30 (54.40) 65.94 (56.46)
EDA, Gaze 59.74 (54.60) 60.80 (53.98) 58.28 (49.39) 57.55 (48.64) 54.12 (47.70) 63.88 (53.86) 56.23 (49.43) 62.29 (55.87) 63.95 (56.89)
EEG, Gaze 60.63 (53.31) 60.01 (53.18) 57.74 (48.44) 57.57 (49.15) 58.90 (49.53) 61.41 (48.68) 57.50 (49.98) 60.22 (52.86) 60.90 (52.04)
ECG, EDA, EEG 58.34 (51.06) 59.96 (52.88) 52.00 (45.33) 50.72 (44.10) 53.38 (47.55) 60.54 (48.45) 53.48 (46.50) 61.66 (54.38) 65.48 (55.72)
ECG, EDA, Gaze 59.79 (52.34) 60.47 (54.12) 57.20 (47.83) 56.51 (47.48) 58.71 (49.91) 63.02 (51.48) 58.05 (51.30) 61.04 (54.06) 62.08 (55.10)
ECG, EEG, Gaze 58.54 (51.16) 58.37 (52.30) 57.63 (49.93) 57.74 (50.91) 56.78 (48.50) 60.34 (47.60) 54.73 (47.46) 59.99 (53.31) 60.97 (51.51)
EDA, EEG, Gaze 61.90 (54.95) 61.55 (55.55) 56.63 (48.09) 56.38 (48.92) 52.31 (45.70) 62.25 (50.91) 58.42 (50.96) 61.95 (55.68) 63.65 (56.13)
All modalities 57.95 (51.58) 63.91 (56.59) 56.35 (50.52) 54.72 (49.67) 54.89 (48.94) 62.73 (51.58) 58.63 (50.49) 63.75 (56.95) 61.38 (55.34)
VI-B Evaluation Protocol and Implementation Details
For all the benchmarks, we use binary (‘high’ versus ‘low’) cognitive load classification. To calculate binary labels, reported scores of less than 5 were considered ‘low’ cognitive load, and scores of 5 or more were considered ‘high’ cognitive load. We use 10-fold and Leave-One-Subject-Out (LOSO) cross-validation evaluation schemes to indicate the performance of our models on previously unseen data. Accuracy and F1 score were used as evaluation metrics. For all the machine learning algorithms except LGBM and XGBoost, we use the Scikit-learn [^8] https://scikit-learn.org/stable/ library. All the parameters of the machine learning algorithms were optimized by performing an exhaustive search.Our deep learning pipelines were implemented using the TensorFlow framework. For the CNN, we used AdaDelta with a decay rate of 0.95, a learning rate of 5e-3, and focal loss (alpha of 4.0 and gamma of 2.0). We empirically selected a batch size of 256 for 100 epochs for all experiments. We implemented our pipelines using an Intel Core i7-9700 CPU and an NVIDIA GeForce RTX 2080 Ti GPU.
@@ VI-C Results
In this evaluation scheme, we observe that the CNN generally has the best (accuracy = 80.30, F1 = 78.34 with ECG, EDA, and GAZE) and the most consistent performance for both accuracy and F1. This is followed by the RF classifier (accuracy = 75.18, F1 = 74.45 with EDA, EEG, and Gaze), which in some cases, outperforms the CNN (e.g., for EDA and/or Gaze). In terms of individual modalities, it can be seen the ECG is the most important, with which the uni-modal classification accuracy and F1 are 78.45 and 76.41, respectively. In the bi-modal setup, we observe that ECG and EDA show the highest results (accuracy = 79.72, F1 = 77.53), followed by ECG and Gaze (accuracy = 78.41, F1 = 76.54). The highest results for tri-modal learning are achieved with ECG, EDA, and Gaze (accuracy = 80.30, F1 = 78.34), closely followed by ECG, EDA, and EEG (accuracy = 80.10, F1 = 78.49). Interestingly, when compared to tri-modal learning results, the performance of the classifiers decreases when all four modalities are used.
@@ VI-C2 LOSO Cross Validation
As the leave-one-subject-out cross-validation evaluation scheme is more challenging than the 10-fold scheme, the performance of classical machine learning and deep learning models decrease in comparison to the 10-fold evaluation scheme. Consistent with our observations for 10-fold evaluation, we observe that for LOSO, CNN has the best performance (accuracy = 65.94, F1 = 56.46), followed by the RF, XGBoost, and LGBM classifier. In some cases, the RF classifier (EEG & Gaze, and ECG, EDA & Gaze) and the LGBM classifier (all modalities) outperform the CNN. In terms of individual modalities, we observe that EEG is the most important modality and yields the best results with accuracy and F1 of 65.21 and 52.15, respectively. Following the same trend, for bi-modal learning, the best results are achieved using EDA and EEG (accuracy = 65.94, F1 = 56.46) followed by ECG and EEG (accuracy = 65.22, F1 = 51.00). The best results for tri-modal learning are achieved with ECG, EDA, and EEG (accuracy = 65.48, F1 = 55.72). When all four modalities (ECG, EDA, EEG, and Gaze) are combined, the best results are achieved with LGBM (accuracy = 63.91, F1 = 56.59).
@@ VI-D Inference Time
As mentioned earlier, in this experiment, the ground truth labels were recorded at 10-second intervals to facilitate the collection of the frequent ground truth labels. This allows for better estimation of cognitive load scores by machine learning models. To meet the real-time inference requirements, the model constraint would be to generate the output prediction in less time than the ground truth collection interval. For classical machine learning models, the inference time would depend on the time to compute handcrafted features and the time to make a prediction. In contrast, for the deep learning model, as there is no computation of handcrafted features, the inference time would depend on the time taken to make a prediction. So, we extract low complexity features for classical machine learning models [45]. Given the number of handcrafted features from four modalities, the average time to compute these features for one sample was approximately 1 second. We then evaluated and present the average inference time of each classical machine learning model and deep learning model in Table VI. The table shows that RF, XGBoost, and CNN models take significantly more time than the other methods; however, the inference time is still under 1 second for all the algorithms that satisfy the real-time inference requirements.

### v2

@@ VI-B Evaluation Protocol and Implementation Details
GB LGBM LDA LR MLP RF SVM XGBoost CNN Transformer
ECG 66.58 (65.49) 66.71 (66.27) 60.49 (57.62) 60.73 (58.21) 62.19 (61.07) 68.28 (67.38) 64.20 (63.83) 66.68 (65.26) 78.45 (76.41) 77.40(66.32)
EDA 68.34 (67.39) 70.18 (69.64) 63.52 (61.65) 63.48 (61.53) 60.56 (59.03) 71.98 (70.93) 65.59 (65.00) 69.33 (68.06) 68.01 (66.21) 74.05(52.84)
EEG 69.19 (68.36) 69.94 (69.42) 64.30 (63.02) 64.84 (63.62) 62.53 (61.94) 71.85 (71.02) 69.09 (68.67) 69.91 (69.00) 74.10 (72.08) 77.37(71.84)
Gaze 66.85 (66.01) 67.97 (67.45) 62.39 (60.11) 62.50 (60.35) 61.54 (59.61) 69.70 (68.66) 65.83 (65.51) 67.80 (66.72) 66.78 (64.23) 74.66(64.88)
ECG, EDA 70.08 (69.31) 71.98 (71.56) 64.37 (62.89) 64.67 (63.20) 59.95 (58.15) 73.24 (72.51) 69.30 (68.83) 70.96 (70.15) 79.72 (77.53) 80.85(73.55)
ECG, EEG 70.76 (69.99) 70.18 (69.67) 65.66 (64.59) 65.93 (64.96) 63.38 (62.56) 72.53 (71.77) 68.75 (68.33) 69.30 (68.32) 77.66 (75.77) 81.85(75.60)
ECG, Gaze 68.96 (68.11) 70.38 (69.92) 63.96 (62.71) 64.13 (62.92) 63.86 (62.67) 71.00 (70.06) 67.80 (67.33) 70.59 (69.71) 78.41 (76.54) 81.00(74.50)
EDA, EEG 72.19 (71.39) 72.09 (71.63) 65.35 (64.21) 65.28 (64.22) 60.96 (59.17) 73.61 (72.85) 68.86 (68.40) 72.49 (71.65) 75.30 (73.45) 82.41(76.76)
EDA, Gaze 71.34 (70.63) 72.32 (71.87) 65.45 (64.12) 65.73 (64.43) 57.67 (54.13) 74.40 (73.65) 68.85 (68.34) 72.63 (71.83) 70.64 (68.23) 81.85(74.99)
EEG, Gaze 71.64 (70.92) 71.71 (71.21) 66.98 (66.08) 67.43 (66.48) 66.30 (64.85) 73.65 (72.76) 70.79 (70.40) 71.40 (70.59) 75.50 (73.13) 82.59(76.79)
ECG, EDA, EEG 71.10 (70.42) 72.90 (72.57) 66.13 (65.07) 65.73 (64.74) 61.71 (60.51) 73.65 (72.97) 69.87 (69.43) 72.53 (71.72) 80.10 (78.49) 84.12(80.08)
ECG, EDA, Gaze 72.12 (71.45) 73.24 (72.86) 67.12 (66.04) 66.61 (65.60) 62.39 (60.10) 74.53 (73.85) 70.96 (70.49) 73.27 (72.54) 80.30 (78.34) 83.43(78.66)
ECG, EEG, Gaze 71.51 (70.82) 72.80 (72.34) 67.56 (66.74) 68.51 (67.67) 66.54 (65.14) 74.09 (73.26) 71.13 (70.80) 71.64 (70.85) 78.26 (76.56) 83.63(78.39)
EDA, EEG, Gaze 73.58 (72.96) 73.58 (73.13) 67.97 (67.13) 67.87 (67.03) 61.51 (59.51) 75.18 (74.45) 71.71 (71.26) 74.50 (73.78) 75.87 (74.08) 84.93(78.18)
All modalities 73.38 (72.81) 74.26 (73.83) 68.24 (67.47) 68.38 (67.59) 62.36 (60.66) 75.14 (74.47) 72.05 (71.59) 74.26 (73.60) 79.65 (77.94) 85.58(81.18)
@@
GB LGBM LDA LR MLP RF SVM XGBoost CNN Transformer
ECG 53.23 (45.21) 42.00 (33.06) 55.60 (43.79) 43.34 (35.22) 43.15 (39.08) 55.60 (44.16) 40.26 (33.09) 40.00 (28.79) 63.28 (51.00) 66.15(54.20)
EDA 58.67 (50.86) 55.20 (51.03) 60.15 (51.23) 50.81 (47.56) 48.90 (43.91) 59.36 (49.50) 50.46 (47.42) 51.52 (47.67) 64.08 (54.80) 63.73(54.43)
EEG 54.71 (47.06) 53.35 (46.40) 49.76 (43.83) 52.38 (44.16) 54.46 (48.94) 59.17 (46.57) 53.23 (47.03) 56.54 (48.95) 65.21 (52.15) 67.77(57.03)
Gaze 54.61 (47.35) 54.79 (46.41) 56.72 (42.49) 52.79 (43.97) 53.08 (44.76) 58.80 (48.83) 51.63 (44.05) 56.93 (46.53) 60.23 (52.92) 65.87(51.73)
ECG, EDA 58.66 (50.54) 45.88 (38.04) 58.37 (47.25) 51.87 (48.68) 47.60 (44.53) 62.23 (51.60) 50.58 (47.58) 47.50 (43.31) 64.48 (55.38) 68.99(62.32)
ECG, EEG 55.10 (48.49) 53.54 (46.77) 50.27 (44.37) 51.97 (44.11) 52.17 (47.11) 56.32 (43.89) 49.29 (43.36) 56.97 (50.15) 65.22 (51.00) 68.74(60.72)
ECG, Gaze 54.29 (49.07) 56.62 (50.75) 57.21 (47.22) 57.22 (47.42) 53.65 (48.40) 57.66 (47.21) 54.51 (47.36) 56.59 (49.79) 59.27 (50.81) 68.02(59.82)
EDA, EEG 56.53 (49.60) 57.91 (52.19) 51.51 (44.89) 51.27 (44.51) 52.80 (46.45) 60.33 (48.58) 53.19 (47.38) 62.30 (54.40) 65.94 (56.46) 70.45(65.00)
EDA, Gaze 59.74 (54.60) 60.80 (53.98) 58.28 (49.39) 57.55 (48.64) 54.12 (47.70) 63.88 (53.86) 56.23 (49.43) 62.29 (55.87) 63.95 (56.89) 69.59(63.20)
EEG, Gaze 60.63 (53.31) 60.01 (53.18) 57.74 (48.44) 57.57 (49.15) 58.90 (49.53) 61.41 (48.68) 57.50 (49.98) 60.22 (52.86) 60.90 (52.04) 69.22(59.51)
ECG, EDA, EEG 58.34 (51.06) 59.96 (52.88) 52.00 (45.33) 50.72 (44.10) 53.38 (47.55) 60.54 (48.45) 53.48 (46.50) 61.66 (54.38) 65.48 (55.72) 70.90(66.84)
ECG, EDA, Gaze 59.79 (52.34) 60.47 (54.12) 57.20 (47.83) 56.51 (47.48) 58.71 (49.91) 63.02 (51.48) 58.05 (51.30) 61.04 (54.06) 62.08 (55.10) 70.61(65.91)
ECG, EEG, Gaze 58.54 (51.16) 58.37 (52.30) 57.63 (49.93) 57.74 (50.91) 56.78 (48.50) 60.34 (47.60) 54.73 (47.46) 59.99 (53.31) 60.97 (51.51) 69.27(62.20)
EDA, EEG, Gaze 61.90 (54.95) 61.55 (55.55) 56.63 (48.09) 56.38 (48.92) 52.31 (45.70) 62.25 (50.91) 58.42 (50.96) 61.95 (55.68) 63.65 (56.13) 72.15(68.50)
All modalities 57.95 (51.58) 63.91 (56.59) 56.35 (50.52) 54.72 (49.67) 54.89 (48.94) 62.73 (51.58) 58.63 (50.49) 63.75 (56.95) 61.38 (55.34) 72.70(69.46)
@@ VI-C Results
In this evaluation scheme, we observe that the Transformer generally has the best performance with an accuracy of 85.58 and an F1 of 81.18 using all the modalities, as well as the most consistent performance for both accuracy and F1. This is followed by the CNN which achieves an accuracy of 80.30 and F1 of 78.34 with ECG, EDA, and Gaze. The CNN, outperforms the Transformer when using ECG. In terms of individual modalities, it can be seen the ECG is the most important, with accuracy and F1 scores of 78.45 and 76.41, respectively, for the CNN. In the bi-modal setup, we observe that EEG and Gaze show the highest results with an accuracy of 82.59 and an F1 of 76.79. This is followed by EDA and EEG which obtain an accuracy of 82.41 and an F1 of 76.76. The highest results for tri-modal learning are achieved with ECG, EDA, and EEG (accuracy = 84.12, F1 = 80.08), closely followed by EDA, EEG, and Gaze (accuracy = 84.93, F1 = 78.18).
@@ VI-C2 LOSO Cross Validation
We observe that in the LOSO setup, the deep learning methods show higher improvements with respect to the classical machine learning method, compared to the 10-fold scheme. Consistent with our observations for 10-fold evaluation, we observe that for LOSO, the Transformer has the best performance with an accuracy of 72.70 and an F1 of 69.46, followed by the CNN and RF methods. In one case (EDA) the CNN classifier outperforms the Transformer. In some cases, the RF classifier (EEG & Gaze, and ECG, EDA & Gaze) and the LGBM classifier (all modalities) outperform the CNN. In terms of individual modalities, we observe that EEG is the most important modality and yields the best results with accuracy and F1 scores of 67.77 and 57.03, respectively. Following the same trend, for bi-modal learning, the best results are achieved using EDA and EEG (accuracy = 70.45, F1 = 65.00) followed by EDA and Gaze (accuracy = 69.59, F1 = 63.20). The best results for tri-modal learning are achieved with EDA, EEG and Gaze (accuracy = 72.15, F1 = 68.50).
@@ VI-D Inference Time
As mentioned earlier, in this experiment, the ground truth labels were recorded at 10-second intervals to facilitate the collection of the frequent ground truth labels. This allows for better estimation of cognitive load scores by machine learning models. To meet the real-time inference requirements, the model constraint would be to generate the output prediction in less time than the ground truth collection interval. For classical machine learning models, the inference time would depend on the time to compute handcrafted features and the time to make a prediction. In contrast, for the deep learning model, as there is no computation of handcrafted features, the inference time would depend on the time taken to make a prediction. So, we extract low complexity features for classical machine learning models [62]. Given the number of handcrafted features from four modalities, the average time to compute these features for one sample was approximately 1 second. We then evaluated and present the average inference time of each classical machine learning model and deep learning model in Table VI. The table shows that RF, XGBoost, CNN and Transformer models take significantly more time than the other methods; however, the inference time is still under 1 second for all the algorithms that satisfy the real-time inference requirements given that each segment is 10 seconds long.